OpenStudy (anonymous):

Match each to the property it best represents. 3/4 = 3/4 If a + 9 = 20, then 20 = a + 9. If 10 = 4x and 4x = c, then 10 = c. -18 + 4 is a unique real number -14. Answer A. Transitive Property of Equality B. Closure of Real Numbers C. Reflexive Property of Equality D. Symmetric Property of Equality

OpenStudy (anonymous):

3/4 = 3/4 -- reflexive property a + 9 = 20, then 20 = a + 9 -- symmetric property 10 = 4x, 4x = c, then 10 = c -- transitive property -18 + 4 is a unique real number -- closure of real numbers

OpenStudy (anonymous):

thank you c: @kelliegirl33

OpenStudy (anonymous):

your quite welcome....everyone needs some help some time or another :)

OpenStudy (anonymous):

yes @kelliegirl33 especially when its math ._. lol

OpenStudy (anonymous):

you are not the only one that has trouble in math....believe me....there are quite a few people that have trouble with math. Join the club....lol